# Sharp Thresholds for Temporal Motifs and Doubling Time in Random Temporal Graphs

### v1

@@ 8.2 Fixed Motifs
Now, while we cannot apply Lemma˜4.5 to the discrete case directly, we are able to exploit the relationship between the two models in order to obtain an analogous result. In particular, we know from Lemma˜4.3 that an occurrence of a $\beta$-temporal motif in the $T$ discretization of $\mathcal{G}$ a sample from $\Gamma_{[n]}(\psi)$ is at most as probable as the occurrence of a $\frac{\beta}{T}$-temporal motif in $\mathcal{G}$. We therefore obtain the following, as an immediate corollary of Lemma˜8.16.
@@
Let $H=(V_{H},E_{H})$ be a fixed static graph, $P=(E_{H},\prec)$ a partial order on its edges, $\beta:\mathbb{N}\to\mathbb{N}$, $T:\mathbb{N}\to\mathbb{N}$ and $\mathcal{G}=(V,E,\lambda)$ a random sample from $\Gamma_{[n]}(\psi,T)$. Then for $\frac{\beta(n)}{T(n)}=o(n^{-\rho_{H}})$ $\mathcal{G}$ does not contain $(H,P)$ as a $\delta$-temporal motif with high probability.
Together these provide the lower bounds on the existence threshold presented in Theorem˜3.1 and Theorem˜3.2 respectively. The upper bounds are slightly more complex to obtain. Our basic strategy is an application of the second moment method using our bounds on the variance obtained in Lemma˜4.6. Unfortunately, due to the obstacles brought about by dependence discussed previously, Lemma˜4.6 only holds when $\psi$ is a Bernoulli random variable. The naive application of the second moment, then, yields the following pair of results (where we defer the near identical proof for the discrete version) which are both considerably weaker than the more general statements we require for Theorem˜3.1 and Theorem˜3.2.
@@
Let $H=(V_{H},E_{H})$ be a fixed static graph, $P=(E_{H},\prec)$ a partial order on its edges, and $\mathcal{G}=(V,E,\lambda)$ a random sample from $\Gamma_{[n]}(\psi)$, such that $\psi$ is a Bernoulli distribution, then for $\delta=\omega(n^{-\rho_{H}})$ $\mathcal{G}$ contains $(H,P)$ as a $\delta$-temporal motif with high probability.
Proof 8.20.
@@
Lemma 8.21.
Let $H=(V_{H},E_{H})$ be a fixed static graph, $P=(E_{H},\prec)$ a partial order on its edges, and $\mathcal{G}=(V,E,\lambda)$ a random sample from $\Gamma_{[n]}(\psi,T)$, such that $\psi$ is a Bernoulli distribution, then for $\frac{\beta}{T}=\omega(n^{\frac{-|V_{H}|}{|E_{H}|-1}})$ $\mathcal{G}$ contains $(H,P)$ as a $\delta$-temporal motif with high probability.
Despite this weakness, however, we are still able to extract the results we need. The basic observation is as follows, for any valid choice of $\psi$, there exists a Bernoulli random variable $\sigma$ stochastically dominated by $\psi$. However, we also know from Lemma˜4.3 that the number of copies of our temporal motif in a sample from $\Gamma_{[n]}(\psi)$ stochastically dominates the number of copies in a sample from $\Gamma_{[n]}(\sigma)$. Together with Lemma˜8.19 and Lemma˜8.21, this is sufficient to obtain the following two results.
@@
Proof 8.23.
Since $\psi$ is a discrete non-negative distribution which is positive with positive probability, there must exist a Bernoulli random variable $\sigma$ such that $\psi$ stochastically dominates $\sigma$. Furthermore, it follows from Lemma˜8.19, that $(H,P)$ exists in $\mathcal{I}=(V,E,\lambda^{\prime})$ with high probability, where $\mathcal{I}$ is sampled from $\Gamma_{[n]}(\sigma)$. From Lemma˜4.3, it thus follows that $(H,P)$ exists within $\mathcal{G}$ with high probability.
@@
Let $H=(V_{H},E_{H})$ be a fixed static graph, $P=(E_{H},\prec)$ a partial order on its edges, and $\mathcal{G}=(V,E,\lambda)$ a random sample from $\Gamma_{[n]}(\psi,T)$, then for $\frac{\beta}{T}=\omega(n^{\frac{-|V_{H}|}{|E_{H}|-1}})$ $\mathcal{G}$ contains $(H,P)$ as a $\delta$-temporal motif with high probability.
Theorem˜3.1 and Theorem˜3.2 then follow immediately from Lemma˜8.16 and Lemma˜8.22, and Corollary˜8.18 and Lemma˜8.24, respectively.
@@ 8.3 Large Cliques
Finally, we shall consider the case where the temporal motifs grow with the size of the temporal graph but $\delta$ the size of the window remains constant. For the sake of brevity, we only consider the continuous model, however one should be able to derive very similar results in the discrete case via the same method. Similarly, for comparison with previous results, we shall focus on the size of the largest motif with a clique footprint and no restriction on the ordering of its edges. We begin with an upper bound on the size of such a motif, which similarly to the case of fixed motifs, we obtain as a consequence of Lemma˜4.5.
@@
Let $\mathcal{G}=(V,E,\lambda)$ be a sample from $\Gamma_{[n]}(\psi)$. Then for any $\epsilon>0$ and $0<\delta<1$, $\mathcal{G}$ does not contain an occurrence of $(H_{k},P^{\emptyset}_{k})$ as a $\delta$-temporal motif with high probability, where
@@
Proof 8.26.
Consider $H$ the clique of size $k$, from Lemma˜4.5,
@@
Unfortunately, the argument used for Lemma˜8.22 is not sufficiently precise for our purposes here, as both the first and second moment of $\psi$ will have a significant impact on the threshold. Instead we make use of Lemma˜4.7, which is better suited for this context as we need to account for the role of $r$ and can absorb the cost of the constant overhead introduced by an additional $\frac{1}{\delta}$ factor. For the case of almost surely bounded random variables, we then obtain the following from Lemma˜4.7 via a straightforward application of the second moment method.
@@
For any $0<\epsilon<\delta<1$, $\psi$ a valid and almost surely bounded label multiplicity distribution and $\mathcal{G}=(V,E,\lambda)$ a sample from $\Gamma_{[n]}(\psi)$. Then, $\mathcal{G}$ contains $(K_{k},P^{\emptyset}_{k})$ as a $\delta$-temporal motif with high probability, where
@@
For some value of $\alpha$ that we will fix later, we shall take $\sigma$ to be the the distribution obtained by truncating $\psi$ to the interval $[0,c]$. We immediately obtain all of the following,
@@
It, then, follows from Lemma˜8.27, that for any choice of $\beta>0$, with high probability, a sample from $\Gamma_{[n]}(\sigma)$ contains $(K_{k^{\prime}},P^{\emptyset}_{k^{\prime}})$ as a $\delta$-temporal motif, where
@@
Theorem˜3.4 then follows immediately from Lemma˜8.25 and Lemma˜8.30.
@@ 9 Doubling Times
In this section, we outline our proof of Theorem˜3.5 characterizing the doubling time of $\Gamma_{[n]}(\psi)$ whenever $\psi$ is a degenerate random variable. The proof largely consists of obtaining probabilistic bounds on the doubling time of small individual sets followed by a more careful analysis of the particularly early and late participants. This approach is inspired by [SharpThresholds], although as we deal with relatively large sets of vertices rather than individual vertices, we have to put in a lot a lot of extra-leg work in the process.
For the sake of mathematical convenience, we will first establish the results for a slightly different random temporal graph model and then show how they can be extended to our original setting. The model in question is a different discretization to the one we have considered thus far: the so-called “order discretization”. In constructing the order discretization of $\mathcal{G}$, we strip away all information about the precise timing of edge labels and instead only retain the relative order between them. The order discretization of $\mathcal{G}$ contains only a single edge at any given time and the edge at time $k$ corresponds to the $k$th edge to appear in $\mathcal{G}$. Formally, it is defined as follows. See 5.1 Our main strategy for obtaining bounds on the doubling time of any specific set is to bound the time required for reachability balls to cross a sequence of cuts in the underlying graph. As a consequence, we will need to make heavy use of the following result bound on the proportion of labels from any cut of the underlying graph sampled in the order discretization before a given time. While in the general case, it is not difficult to show that any fixed cut, determined before any sampling occurs, has this property, we need to select the cuts for consideration based on the growth of the ball so far. This introduces a great deal of dependence which we circumvent via the following result.
@@
$\mathbb{P}\left[\max_{S\subseteq V}\frac{\sum_{e\in S\times V\setminus S}\lambda^{\prime}_{[0,\alpha n\log{n}]}(e)}{(r|S|(n-|S|))}>\beta\right]=o(n^{-1}).$
@@ 9.1 Upper Bounds
For ease of notation we will denote $D=\textbf{Double}^{+}(\mathcal{G})$ and $\mathcal{H}=\mathcal{R}(\mathcal{G},\textbf{Double}^{+}(\mathcal{G})$. Now for the sake of contradiction, assume that there exists a set $S$ with $|S|\leq\frac{n}{4}$ such that $\text{Double}_{\mathcal{H}}(S)\geq D$. Consider that for any two vertices $u$ and $v$ if there exists a path from $u$ to $v$ in $\mathcal{G}$ arriving before time $D$ there must exist a path from $v$ to $u$ in $\mathcal{H}$ arriving by time $D$. Thus, we must have that $T$ the set of vertices outside of $S$ that can reach $S$ in $\mathcal{G}$ before time $D$ has size at most |S|1|S|-1. However, since $S\leq\frac{n}{4}$ we must have that $|S\cup T|\leq\frac{n}{2}$ and so we can take $U\subset V$ such that $|U|=\frac{n}{2}$ and $S\cup T\subseteq U$. Now by definition $V\setminus U$ has doubling time at most $D$ and so there must some vertex from $V\setminus U$ with a path to a member of $S$ in $\mathcal{G}$ arriving by time $D$. This contradicts the definition of $T$ and so we obtain our result.
While stated for arbitrary temporal graphs, in our setting, given the equal probability of sampling both $\mathcal{G}$ and $(\mathcal{R}(\mathcal{G},\textbf{Double}^{+}(\mathcal{G}))$, this turns any upper bound on the doubling time of large sets into an upper bound on the doubling time of small sets. Our second key preliminary result is an upper bound on the time taken for the reachability ball centered on any specific set of vertices to grow to a certain a size. We obtain the result by a carefully constructed coupling between the total time between expansions of the reachability ball and a sum of geometric random variables, but defer the somewhat laborious calculations. Denote by $\mathbf{H}(a,b)=H_{n-a}+H_{b-1}-H_{a-1}-H_{n-b}$, where $H_{a}$ is the $a$th harmonic number with the convention that $H_{0}=0$.
@@
For any $\alpha\geq 0$, let $\mathcal{H}_{\alpha}=\mathcal{G}^{O}_{[\alpha n\log{n},\infty)}$, where $\mathcal{G}=(V,E,\lambda)$ is a sample from $\Gamma_{[n]}(\psi)$ and $\psi$ is a constant random variable. Then for any $S\subseteq V$ chosen with knowledge of $\mathcal{G}^{O}_{[0,\alpha n\log{n}]}$, $|S|<k\leq n$, $\beta>0$ and $0<\gamma<1$ small, we have the following.
@@
With these tools established, we may give our upper bound on the doubling time in $\mathcal{G}^{O}$. See 5.4 The proof of this result consists of largely of applying Lemma˜9.3, along with a concentration inequality and Lemma˜5.3 in order to deal with the case of small sets. More specifically, observe that once all vertices can be reached by at least $\frac{3n}{4}$ other vertices, every set of size at least $\frac{n}{4}$ can reach all vertices. Thus, once this has occurred the reachability ball of all large sets of vertices must have doubled in size. In order to bound this time, we first show via an application of Lemma˜9.3 and Markov’s inequality that almost all vertices can be reached by $\frac{3n}{4}$ other vertices by roughly time $n\log{n}$. From a second application of Lemma 9.3 we find that the few remaining vertices are reached from the already well connected vertices within only a further approximately $n\log{n}$ rounds. The more general result then follows from using Lemma˜5.3 to extend the bound to cover all smaller sets of vertices.
@@
Thus with probability $1-\frac{1}{\log{n}}$ we have at set of at least $n-n^{1-\epsilon+\log{1+\epsilon}}\log{n}$ vertices that can be reached by at least $\frac{3n}{4}$ other vertices which we shall denote by $A$. Trivially, $T$ must occur at the very latest when every member of $V\setminus A$ is reachable by a member of $A$ via a path starting after $T_{1}$. Since $T_{1}=O(n\log{n})$, we know from a second application of Lemma˜9.3 that this occurs at the very latest by $T_{1}+\frac{(1+\epsilon)n\mathbf{H}(n-n^{1-\epsilon+\log{1+\epsilon}}\log{n},n)}{2(1-\gamma)}$ with probability 1o(1)1-o(1). Taking a union bound over the probability of failures we get that,
@@
$\leq\frac{(1+\epsilon)n}{2(1-\gamma)}\left(2(1+\alpha)\log{n}-(1-\alpha^{2})\log{n}+(1+\alpha)\log{n}-(1-\alpha)\log{n}+(1+\alpha)(1-\epsilon+\log{(1+\epsilon)})\log{n}\right)$
$\frac{n\log{n}}{2(1-\gamma)}(1+\epsilon)\left[\left(4-\epsilon+\log{(1+\epsilon)}\right)(1+\alpha)-(2+\alpha)(1-\alpha)\right]$
@@
Thus, we find that with probability 1o(1)1-o(1) that all vertices can be reached by at least $\frac{3n}{4}$ other vertices by time $(1+\beta)n\log{n}$. Therefore, we must have that by time $(1+\beta)n\log{n}$ all sets of size at most $\frac{n}{2}$ and $\frac{n}{4}$ can reach all vertices. Thus we have that with probability 1o(1)1-o(1), $\textbf{Double}^{+}(\mathcal{G})\leq(1+\beta)n\log{n}$ and by the time reversible nature of the distribution that $\textbf{Double}^{+}(\mathcal{R}(\mathcal{G},(1+\beta)n\log{n}))\leq(1+\beta)n\log{n}$ with the same probability. It follows from Lemma˜5.3 that $\textbf{Double}^{-}(\mathcal{G})\leq(1+\beta)n\log{n}$, and so taking a union bound over the failure probability that $\textbf{Double}(\mathcal{G})\leq(1+\beta)n\log{n}$

### v2

@@ 8.2 Fixed Motifs
Now, while we cannot apply Lemma˜4.5 to the discrete case directly, we are able to exploit the relationship between the two models in order to obtain an analogous result. In particular, we know from Lemma˜4.3 that an occurrence of a $\beta$-temporal motif in the $T$ discretization of $\mathcal{G}$, sampled from $\Gamma_{[n]}(\psi)$, is at most as probable as the occurrence of a $\frac{\beta}{T}$-temporal motif in $\mathcal{G}$. We therefore obtain the following, as an immediate corollary of Lemma˜8.18.
@@
Let $H=(V_{H},E_{H})$ be a fixed static graph, $P=(E_{H},\prec)$ a partial order on its edges, $\beta:\mathbb{N}\to\mathbb{N}$, $T:\mathbb{N}\to\mathbb{N}$ and $\mathcal{H}=(V,E,\lambda)$ a random sample from $\Gamma_{[n]}(\psi,T)$. Then for $\frac{\beta(n)}{T(n)}=o(n^{-\rho_{H}})$, $\mathcal{H}$ does not contain $(H,P)$ as a $\beta$-temporal motif with high probability.
Together these provide the lower bounds on the existence threshold presented in Theorem˜3.1 and Theorem˜3.2 respectively. The upper bounds are slightly more complex to obtain. Our basic strategy is an application of the second moment method using our bounds on the variance obtained in Lemma˜4.6. Unfortunately, due to the obstacles brought about by dependence discussed previously, Lemma˜4.6 only holds when $\psi$ is a Bernoulli random variable. The naive application of the second moment, then, yields the following pair of results (where we defer the near identical proof of the discrete version) which are both considerably weaker than the more general statements we require for Theorem˜3.1 and Theorem˜3.2.
@@
Let $H=(V_{H},E_{H})$ be a fixed static graph, $P=(E_{H},\prec)$ a partial order on its edges, and $\mathcal{G}=(V,E,\lambda)$ a random sample from $\Gamma_{[n]}(\psi)$, such that $\psi$ is a Bernoulli distribution, then for $\delta=\omega(n^{-\rho_{H}})$, $\mathcal{G}$ contains $(H,P)$ as a $\delta$-temporal motif with high probability.
Lemma 8.22.
Let $H=(V_{H},E_{H})$ be a fixed static graph, $P=(E_{H},\prec)$ a partial order on its edges, and $\mathcal{H}=(V,E,\lambda)$ a random sample from $\Gamma_{[n]}(\psi,T)$, such that $\psi$ is a Bernoulli distribution, then for $\frac{\beta}{T}=\omega(n^{\frac{-|V_{H}|}{|E_{H}|-1}})$ $\mathcal{H}$ contains $(H,P)$ as a $\delta$-temporal motif with high probability.
Proof 8.23 (Proof of Lemma˜8.21).
@@
Despite this weakness, however, we are still able to extract the results we need. The basic observation is as follows, for any valid choice of $\psi$, there exists a Bernoulli random variable $\sigma$ stochastically dominated by $\psi$. However, we also know from Lemma˜4.3 that the number of copies of our temporal motif in a sample from $\Gamma_{[n]}(\psi)$ stochastically dominates the number of copies in a sample from $\Gamma_{[n]}(\sigma)$. Together with Lemma˜8.21 and Lemma˜8.22, this is sufficient to obtain the following two results.
@@
Proof 8.25.
Since $\psi$ is a discrete non-negative distribution which is positive with positive probability, there must exist a Bernoulli random variable $\sigma$ such that $\psi$ stochastically dominates $\sigma$. Furthermore, it follows from Lemma˜8.21, that $(H,P)$ exists in $\mathcal{I}=(V,E,\lambda^{\prime})$ with high probability, where $\mathcal{I}$ is sampled from $\Gamma_{[n]}(\sigma)$. From Lemma˜4.3, it thus follows that $(H,P)$ exists within $\mathcal{G}$ with high probability.
@@
Let $H=(V_{H},E_{H})$ be a fixed static graph, $P=(E_{H},\prec)$ a partial order on its edges, and $\mathcal{H}=(V,E,\lambda)$ a random sample from $\Gamma_{[n]}(\psi,T)$, then for $\frac{\beta}{T}=\omega(n^{\frac{-|V_{H}|}{|E_{H}|-1}})$, $\mathcal{H}$ contains $(H,P)$ as a $\delta$-temporal motif with high probability.
Theorem˜3.1 and Theorem˜3.2 then follow immediately from Lemma˜8.18 and Lemma˜8.24, and Corollary˜8.20 and Lemma˜8.26, respectively.
@@ 8.3 Large Cliques
Finally, we shall consider the case where the temporal motifs grow with the size of the temporal graph but $\delta$ the size of the window remains constant. For the sake of brevity, we only consider the continuous model. However, one should be able to derive very similar results in the discrete case via the same method. Similarly, for comparison with previous results, we shall focus on the size of the largest motif with a clique footprint and no restriction on the ordering of its edges. We begin with an upper bound on the size of such a motif, which similarly to the case of fixed motifs, we obtain as a consequence of Lemma˜4.5.
@@
Let $\mathcal{G}=(V,E,\lambda)$ be a sample from $\Gamma_{[n]}(\psi)$. Then for any $0<\epsilon<\delta<1$, $\mathcal{G}$ does not contain an occurrence of $(K_{k},P^{\emptyset}_{k})$ as a $\delta$-temporal motif with high probability, where
@@
Proof 8.28.
Consider $H=K_{k}$ the clique of size $k$, from Lemma˜4.5,
@@
Unfortunately, the argument used for Lemma˜8.24 is not sufficiently precise for our purposes here, as both the first and second moment of $\psi$ will have a significant impact on the threshold. Instead we make use of Lemma˜4.7, which is better suited for this context as we need to account for the role of $r$ but can absorb the cost of the constant overhead introduced by an additional $\frac{1}{\delta}$ factor. For the case of almost surely bounded random variables, we then obtain the following from Lemma˜4.7 via a straightforward application of the second moment method.
@@
For any $0<\epsilon<\delta<1$, $\psi$ a valid and almost surely bounded label multiplicity distribution and $\mathcal{G}=(V,E,\lambda)$ a sample from $\Gamma_{[n]}(\psi)$, we have that $\mathcal{G}$ contains $(K_{k},P^{\emptyset}_{k})$ as a $\delta$-temporal motif with high probability, where
@@
For some value of $\alpha$ that we will fix later, we shall take $\sigma$ to be the distribution obtained by truncating $\psi$ to the interval $[0,c]$. We immediately obtain all of the following,
@@
It, then, follows from Lemma˜8.29, that for any choice of $\beta>0$, with high probability, a sample from $\Gamma_{[n]}(\sigma)$ contains $(K_{k^{\prime}},P^{\emptyset}_{k^{\prime}})$ as a $\delta$-temporal motif, where
@@
Theorem˜3.4 then follows immediately from Lemma˜8.27 and Lemma˜8.32.
@@ 9 Doubling Times
In this section, we outline our proof of Theorem˜3.5 characterizing the doubling time of $\Gamma_{[n]}(\psi)$ whenever $\psi$ is a degenerate random variable. The proof largely consists of obtaining probabilistic bounds on the doubling time of small individual sets, followed by a more careful analysis of the particularly early and late participants. This approach is inspired by [SharpThresholds], although as we deal with relatively large sets of vertices rather than individual vertices, we have to put in a lot of extra-leg work in the process.
For the sake of mathematical convenience, we first establish the results for a slightly different random temporal graph model and then show how they can be extended to our original setting. The model in question is a different discretization to the one we have considered thus far: the so-called “order discretization”. In constructing the order discretization of $\mathcal{G}$, we strip away all information about the precise timing of edge labels and instead only retain the relative order between them. The order discretization of $\mathcal{G}$ contains only a single edge at any given time and the edge at time $k$ corresponds to the $k$th edge to appear in $\mathcal{G}$. Formally, it is defined as follows. See 5.1 Our main strategy for obtaining bounds on the doubling time of any specific set is to bound the time required for reachability balls to cross a sequence of cuts in the underlying graph. As a consequence, we will need to make heavy use of the following result bound on the proportion of labels from any cut of the underlying graph sampled in the order discretization before a given time. While in the general case, it is not difficult to show that any fixed cut, determined before any sampling occurs, has this property, we need to select the cuts for consideration based on the growth of the ball so far. This introduces a great deal of dependence which we circumvent via the following result.
@@
$\mathbb{P}\left[\max_{S\subseteq V}\frac{\sum_{e\in S\times V\setminus S}\lambda^{O}_{[0,\alpha n\log{n}]}(e)}{(r|S|(n-|S|))}>\beta\right]=o(n^{-1}).$
@@ 9.1 Upper Bounds
For ease of notation we will denote $D=\textbf{Double}^{+}(\mathcal{G})$ and $\mathcal{H}=\mathcal{R}(\mathcal{G},\textbf{Double}^{+}(\mathcal{G}))$. Now for the sake of contradiction, assume that there exists a set $S$ with $|S|\leq\frac{n}{4}$ such that $\text{Double}_{\mathcal{H}}(S)\geq D$. Consider that for any two vertices $u$ and $v$ if there exists a path from $u$ to $v$ in $\mathcal{G}$ arriving before time $D$ there must exist a path from $v$ to $u$ in $\mathcal{H}$ arriving by time $D$. Thus, we must have that $T$ the set of vertices outside of $S$ that can reach $S$ in $\mathcal{G}$ before time $D$ has size at most |S|1|S|-1. Furthermore, since $S\leq\frac{n}{4}$, we must have that $|S\cup T|\leq\frac{n}{2}$ and so we can take $U\subset V$ such that $|U|=\frac{n}{2}$ and $S\cup T\subseteq U$. Now by definition $V\setminus U$ has doubling time at most $D$. Since $V\setminus U$ has size $\frac{n}{2}$, this implies that every vertex can be reached from $V\setminus U$ via a time-respecting path before time $D$. So, in particular, there must be at least one vertex from $V\setminus U$ with a path to some of $S$ in $\mathcal{G}$ arriving by time $D$. In turn, this implies that some member of $V\setminus U$ can be reached by a member of $S$ before time $D$ in $\mathcal{H}$. This contradicts the definition of $T$ and so we obtain our result.
While stated for arbitrary temporal graphs, given the equal probability of sampling both $\mathcal{G}$ and $\mathcal{R}(\mathcal{G},\textbf{Double}^{+}(\mathcal{G}))$ in our setting, this turns any upper bound on the doubling time of large sets into an upper bound on the doubling time of small sets. Our second key preliminary result is an upper bound on the time taken for the reachability ball centered on any specific set of vertices to grow to a certain size. We obtain the result by a carefully constructed coupling between the total time between expansions of the reachability ball and a sum of geometric random variables, but defer the somewhat laborious calculations. Denote by $\mathbf{H}(a,b)=H_{n-a}+H_{b-1}-H_{a-1}-H_{n-b}$, where $H_{a}$ is the $a$th harmonic number with the convention that $H_{0}=0$.
@@
For any $\alpha\geq 0$, let $\mathcal{H}_{\alpha}=\mathcal{G}^{O}_{[\alpha n\log{n},\infty)}$, where $\mathcal{G}=(V,E,\lambda)$ is a sample from $\Gamma_{[n]}(\psi)$ with $\psi$ a constant random variable. Then for any $S\subseteq V$ chosen with knowledge of $\mathcal{G}^{O}_{[0,\alpha n\log{n}]}$, $|S|<k\leq n$, $\beta>0$ and $0<\gamma<1$ small, we have the following.
@@
With these tools established, we may give our upper bound on the doubling time in $\mathcal{G}^{O}$. See 5.4 The proof of this result consists largely of applying Lemma˜9.3, along with a concentration inequality and Lemma˜5.3 in order to deal with the case of small sets. More specifically, observe that once all vertices can be reached by at least $\frac{3n}{4}$ other vertices, every set of size at least $\frac{n}{4}$ can reach all vertices. Thus, once this has occurred the reachability ball of all large sets of vertices must have doubled in size. In order to bound this time, we first show via an application of Lemma˜9.3 and Markov’s inequality that almost all vertices can be reached by $\frac{3n}{4}$ other vertices by roughly time $n\log{n}$. From a second application of Lemma˜9.3 we find that the few remaining vertices are reached from the already well connected vertices within only approximately a further $n\log{n}$ rounds. The more general result then follows from using Lemma˜5.3 to extend the bound to cover all smaller sets of vertices.
@@
Thus with probability $1-\frac{1}{\log{n}}$ we have a set of at least $n-n^{1-\epsilon+\log{1+\epsilon}}\log{n}$ vertices that can be reached by at least $\frac{3n}{4}$ other vertices which we shall denote by $A$. Trivially, $T$ must occur at the very latest when every member of $V\setminus A$ is reachable by a member of $A$ via a path starting after $T_{1}$. Since $T_{1}=O(n\log{n})$, we know from a second application of Lemma˜9.3 that this occurs at the very latest by $T_{1}+\frac{(1+\epsilon)n\mathbf{H}(n-n^{1-\epsilon+\log{1+\epsilon}}\log{n},n)}{2(1-\gamma)}$ with probability 1o(1)1-o(1). Taking a union bound over the probability of failures we get that with probability 1o(1)1-o(1),
@@
$\leq\frac{(1+\epsilon)n\log{n}}{2(1-\gamma)}\left(\left(4-\epsilon+\log{(1+\epsilon)}+\frac{\log\log{n}}{\log{n}}\right)(1+\alpha)-2(1-\alpha)\right)$
$\frac{(1+\epsilon)n\log{n}}{2(1-\gamma)}\left(2+2\alpha+(1+\alpha)\left(-\epsilon+\log{(1+\epsilon)}+\frac{\log\log{n}}{\log{n}}\right)\right)$
@@
Thus, we find that, with probability 1o(1)1-o(1), all vertices can be reached by at least $\frac{3n}{4}$ other vertices by time $(1+\beta)n\log{n}$. Therefore, we must have that by time $(1+\beta)n\log{n}$ all sets of size at least $\frac{n}{4}$ can reach all vertices. Thus we have that with probability 1o(1)1-o(1), $\textbf{Double}^{+}(\mathcal{G})\leq(1+\beta)n\log{n}$ and by the time reversible nature of the distribution that $\textbf{Double}^{+}(\mathcal{R}(\mathcal{G},(1+\beta)n\log{n}))\leq(1+\beta)n\log{n}$ with the same probability. In which case it follows from Lemma˜5.3 that $\textbf{Double}^{-}(\mathcal{G})\leq(1+\beta)n\log{n}$, and so taking a union bound over the failure probabilities, we find that $\textbf{Double}(\mathcal{G})\leq(1+\beta)n\log{n}$